In an ore containing Uranium, the ratio of U238 to Pb206 nuclei is 3. Calculate the age of the ore, assuming that all the lead present in the ore is the final stable product of U238 . Take the half-life of U238 to 4.5 x 109 years

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Hello Student,
Please find the answer to your question
Number of atoms of 238U initially / Numbarof atoms of 238U finally = 4 / 3 = a / (a - x)
[∵ Initially one part lead is present with three parts Uranium]
∴ t = 2.303 / λ log α / (α – x )= 2.303 x 4. 5 x 109 / 0.693 log 4/3
= 1.868 x 109 years.
Thanks
Navjot Kalra
askIITians Faculty